# RoboEXP: Action-Conditioned Scene Graph via Interactive Exploration for Robotic Manipulation

### v1

@@ I-A Interactive Exploration
I-B Usage of ACSG
Judging Object Existence. A direct application of our ACSG is to determine the presence or absence of specific objects in the current environment. For instance, during the exploitation stage of the scenario to set the dining table, if the spoon is missing, the robot can further seek human assistance.
Object Retrieval. One notable advantage of our ACSG is its ability to capture all actions and their preconditions. Utilizing this information, retrieving any object becomes straightforward by following the graph structure and executing actions in topological order along the paths from the root to the target object node. For example, in the obstruction scenario, the ACSG can provide the sequence of actions required to fetch the tape: 1) removing the condiment blocking the cabinet door, 2) opening the cabinet via the door handle, and 3) retrieving the tape. Such insights are crucial for tasks like cooking.
I-C Decision-Making Module
@@
I-D Action Module
The action module focuses on providing useful action primitives to aid in constructing our ACSG. We have designed seven action primitives: “open the [door]”, “open the [drawer]”, “close the [door]”, “close the [drawer]”, “pick [object] to idle space”, “pick back [object]”, “move wrist camera to [position]”. To fully support autonomous actions, we employ a heuristic-based algorithm leveraging geometric cues.
For the door and drawer relevant primitives, engagement with handles is required. In our implementation, we exploit the handle’s position and geometry to discern its motion type (prismatic or revolute) and motion parameters (motion axis and motion origin). Executing this action involves utilizing the detected handle and its geometry to adeptly open doors or drawers. Upon identifying the specific handle to be operated, our system retrieves the point cloud converted from our voxel-based representation corresponding to that handle from our memory module. Subsequently, we employ Principal Component Analysis (PCA) to determine the principal direction of the handle, aiding in aligning the gripper for optimal engagement. Additionally, understanding the opening direction is pivotal for effectively handling doors or drawers. To ascertain this, we analyze neighboring points and deduce the most common normal as the opening direction. The combined information of the handle direction and the opening direction provides sufficient guidance for our robot arm to grasp the handle and open the prismatic part. However, in the case of a revolute joint, the motion becomes more intricate. Therefore, we further utilize the motion parameters inferred from the geometry to simulate the evolving opening direction based on the revolute joint’s opening process. This well-designed heuristic empowers our system to reliably open drawers or doors in our tabletop setting.
@@ I-D Action Module
II Additional Details of Experiments
II-A GPT-4V Baseline
We have developed the pure GPT-4V baseline, incorporating ground truth action primitives (where humans manually perform the actions), establishing it as a robust benchmark for comparison with our approach. The full prompt of the GPT-4V baseline is illustrated in Fig. 8.
II-B Experiment Settings
@@
To assess the effectiveness of our system, we devised five types of experiments, each encompassing 10 distinct settings. These settings vary in terms of object number, type, and layout, as illustrated in Fig. 10.
II-C Human Intervention
@@
Our RoboEXP system possesses the capability to autonomously adapt to changes in the environment. We employ two types of human interventions to demonstrate these points.

### v2

@@ B.1.2 Memory Module
B.1.3 Decision-Making Module
@@
We adhere to the standard practice for designing prompts, as other papers do using LLM/LMM [83, 89, 18]. In order not to compromise the generalization ability of our system, we consistently use the same prompts across all scenarios and experiments. Our fundamental rule for prompt design is to minimize ambiguity and ensure alignment with our task. In our experiments, the average response time from GPT-4V is about 8 seconds for each question, which is acceptable as GPT-4V is only used in high-level task planning. For low-level motion planning, the use of action primitives allows us to meet the high-frequency requirement without having to constantly querying GPT-4V.
Our ACSG utilizes GPT-4V on every object node to progressively expand the graph. Hence, regardless of how complicated the scene is, each query posed to GPT-4V resides on a local node within our ACSG, essentially addressing the question, “Should I proceed with exploring this object, and if so, how?” As shown in our Matryoshka scenario, RoboEXP performs well in complex scenarios featuring five levels of hierarchical scene graphs and complicated exploration procedures. The commonsense knowledge learned by GPT-4V enables our system to efficiently explore the environment without having to manually design the exploration rules for diverse objects.
B.1.4 Action Module
The action module focuses on providing useful action primitives to aid in constructing our ACSG. We have designed seven action primitives: “open the [door]”, “open the [drawer]”, “close the [door]”, “close the [drawer]”, “pick [object] to idle space”, “pick back [object]”, “move wrist camera to [position]”. To fully support autonomous actions, we employ a heuristic-based algorithm leveraging geometric cues. The input to each action primitive is an object node of our ACSG, from which we can extract all necessary semantic, state, and low-level geometry information of the object. The integrated information can help us determine the specific grasping pose and path planning for opening and picking actions, which generalize to different instances in various positions and poses.
For opening action primitives related to doors or drawers, engagement with handles is required. In our implementation, we exploit the handle’s position and geometry to discern its motion type (prismatic or revolute) and motion parameters (motion axis and motion origin). Executing this action involves utilizing the detected handle and its geometry to adeptly open doors or drawers. Upon identifying the specific handle to be operated, our system retrieves the point cloud converted from our voxel-based representation corresponding to that handle from our memory module. Subsequently, we employ Principal Component Analysis (PCA) to determine the principal direction of the handle, aiding in aligning the gripper for optimal engagement. Additionally, understanding the opening direction is pivotal for effectively handling doors or drawers. To ascertain this, we analyze neighboring points and deduce the most common normal as the opening direction. The combined information of the handle direction and the opening direction provides sufficient guidance for our robot arm to grasp the handle and open the prismatic part. However, in the case of a revolute joint, the motion becomes more intricate. Therefore, we further utilize the motion parameters inferred from the geometry to simulate the evolving opening direction based on the revolute joint’s opening process. This well-designed heuristic empowers our system to reliably open drawers or doors in our tabletop setting.
@@ B.1.4 Action Module
B.2 Other Design in Interactive Exploration
One desiderata for robot exploration is the ability to handle scenarios that necessitate multi-step or recursive reasoning. An example of this is the Matryoshka doll case, which cannot be addressed using previous one-step LLM-based code generation approaches [19, 18]. In contrast, our modular design allows agents to dynamically plan and adapt in a closed-loop manner, enabling continuous LLM-based exploration based on environmental feedback.
B.3 Usage of ACSG
Judging Object Existence. A direct application of our ACSG is to determine the presence or absence of specific objects in the current environment. For instance, during the exploitation stage of the scenario (App. D) to set the dining table, if the spoon is missing, the robot can further seek human assistance.
Object Retrieval. One notable advantage of our ACSG is its ability to capture all actions and their preconditions. Utilizing this information, retrieving any object becomes straightforward by following the graph structure and executing actions in topological order along the paths from the root to the target object node. For example, in the obstruction scenario (App. D), the ACSG can provide the sequence of actions required to fetch the tape: 1) removing the condiment blocking the cabinet door, 2) opening the cabinet via the door handle, and 3) retrieving the tape. Such insights are crucial for tasks like cooking.
Appendix C Additional Details of Experiments
We conduct experiments in different settings to validate the effectiveness of the model. We provide additional experiments and results, including those with different lighting conditions and backgrounds, using different LMMs, intervention experiments, and several more room-level scenarios.
C.1 Experiment Settings
@@
Our experimental setup encompasses a diverse range of objects, as illustrated in Fig. 8. To assess the effectiveness of our system, we devised five types of experiments for the main quantitative results, each encompassing 10 distinct settings. These settings vary in terms of object number, type, and layout, as illustrated in Fig. 13.
C.2 Evaluation
The details of five metrics employed for evaluation are as follows:
C.3 Baselines
GPT-4V Baseline: We employ the pure GPT-4V as our baseline model along with the chain-of-thoughts (CoT) to enhance its capabilities, as outlined in a method similar to that proposed by Hu et al. [19]. The full prompt of the GPT-4V baseline is illustrated in Fig. 9.
Heuristic Baseline: We also design two heuristic baselines: i) Heuristic-Open: This method opens all handles; ii) Heuristic-Full: In addition to opening all handles, this method also picks up all movable objects to reveal what’s underneath them.
Random Baseline: For the Random baseline, an action is chosen randomly for each object, including the option to take no action. Similar to the GPT-4V baseline, we make the heuristic and random baselines stronger by assuming that all actions are successful. We conducted the same set of experiments as in our original paper (5 task scenarios with 10 different settings for each scenario) for each of the baselines. All methods use the same perception module as our RoboEXP to understand the influence of different decision strategies.
Table 2: Quantitative Results on Different Tasks. We compare the performance of baselines and our system across various tasks. Our system outperforms the baselines across nearly all metrics.
[…]
Drawer-Only
Random 0$\pm$0.0 20$\pm$13.3 10$\pm$10 85$\pm$7.6 6.6$\pm$0.73
Heuristic-Open 90$\pm$10.0 97$\pm$3.3 100$\pm$0.0 0$\pm$0.0 0.2$\pm$0.20
Heuristic-Full 0$\pm$0.0 97$\pm$3.3 100$\pm$0.0 0$\pm$0.0 3.8$\pm$0.47
GPT-4V 20$\pm$13.3 83$\pm$11.0 60$\pm$16.3 15$\pm$7.6 2.8$\pm$1.04
Ours 90$\pm$10.0 97$\pm$3.3 100$\pm$0.0 0$\pm$0.0 0.2$\pm$0.20
Door-Only
Random 0$\pm$0.0 30$\pm$15.3 0$\pm$0.0 85$\pm$7.64 6.3$\pm$1.01
Heuristic-Open 90$\pm$10.0 100$\pm$0.0 100$\pm$0.0 0$\pm$0.0 0.1$\pm$0.10
Heuristic-Full 0$\pm$0.0 100$\pm$0.0 100$\pm$0.0 0$\pm$0.0 3.8$\pm$0.47
GPT-4V 30$\pm$15.2 50$\pm$16.7 80$\pm$13.3 40$\pm$14.5 4.4$\pm$1.52
Ours 90$\pm$10.0 100$\pm$0.0 100$\pm$0.0 0$\pm$0.0 0.1$\pm$0.10
Drawer-Door
Random 0$\pm$0.0 4$\pm$4.0 0$\pm$0.0 85$\pm$6.7 12.8$\pm$0.80
Heuristic-Open 70$\pm$15.3 91$\pm$4.7 100$\pm$0.0 0$\pm$0.0 0.5$\pm$0.27
Heuristic-Full 0$\pm$0.0 91$\pm$4.7 100$\pm$0.0 0$\pm$0.0 6.1$\pm$0.71
GPT-4V 10$\pm$10.0 62$\pm$10.7 70$\pm$15.3 25$\pm$6.6 5.6$\pm$1.46
Ours 70$\pm$15.3 91$\pm$4.7 100$\pm$0.0 0$\pm$0.0 0.5$\pm$0.27
Recursive
Random 10$\pm$10.0 33$\pm$12.9 60$\pm$16.3 70$\pm$12.8 8.8$\pm$1.91
Heuristic-Open 0$\pm$0.0 0$\pm$0.0 100$\pm$0.0 100$\pm$0.0 10$\pm$1.69
Heuristic-Full 40$\pm$16.3 90$\pm$10.0 100$\pm$0.0 5$\pm$5.0 1.8$\pm$0.63
GPT-4V 0$\pm$0.0 20$\pm$13.3 70$\pm$15.3 63$\pm$15.3 8.8$\pm$2.06
Ours 70$\pm$15.3 80$\pm$11.7 100$\pm$0.0 15$\pm$8.9 2.1$\pm$1.49
Occlusion
Random 0$\pm$0.0 0$\pm$0.0 0$\pm$0.0 100$\pm$0.0 9.8$\pm$0.53
Heuristic-Open 0$\pm$0.0 22$\pm$11.7 0$\pm$0.0 75$\pm$8.3 8.6$\pm$0.72
Heuristic-Full 0$\pm$0.0 22$\pm$11.7 0$\pm$0.0 75$\pm$8.3 9.4$\pm$0.58
GPT-4V 0$\pm$0.0 17$\pm$11.4 10$\pm$10.0 85$\pm$7.6 7.3$\pm$0.97
Ours 50$\pm$16.7 67$\pm$14.9 70$\pm$15.3 30$\pm$15.3 2.5$\pm$1.15
Table 3: We report the average number of actions for settings in which all methods find all objects, explore all hidden spaces, and recover the states before exploration. In the recursive and occlusion scenarios, some baseline methods do not have experiments that meet these requirements. Our RoboEXP consistently requires fewer actions compared to the baselines and is more aligned with the ground truth of human action choices.
Method Drawer-Only Door-Only Drawer-Door Recursive Occlusion
GT (Human) 4.0 4.0 8.0 4.6 6.0
Heuristic-Full 7.0 8.0 12.0 6.0 -
GPT-4V 6.0 4.4 8.0 - -
Ours 4.0 4.0 8.0 4.6 6.3
Figure 10: Experiments on Extreme Illumination and Random Background. We conduct experiments in four scenarios with varying lighting conditions and random backgrounds. The reported numbers are averages over four scenarios for each condition. Our system performs well across all conditions. (OR refers to our Object Recovery metric).
C.3.1 Comparison to Heuristic & Random Baseline
We present all results in the above table in Tab. 2, which shows that our RoboEXP system is more capable and efficient in nearly all scenarios. The random baseline easily fails due to improper action selections. The Heuristic-Open baseline performs well in scenarios involving only open-action judgments but fails when pick actions are required. The Heuristic-Full baseline can provide better object recovery; however, its redundant pick actions create a scene graph with too many redundant nodes, resulting in higher Graph Edit Distance in most scenarios. In complicated scenarios like object-object occlusion, the heuristic baselines completely fail, as they do not reason about object-object relationships as we do using the action verifier in our system.
There are two main reasons why heuristic-based baselines are not ideal: i) Defining all potential heuristics to cover every scenario is nearly impossible. Specifying and enumerating all semantic categories is tedious, and some size-based heuristics can also fail (e.g., a threshold that is too small may filter out objects that should be picked up, such as clothes on a Kindle, while a threshold that is too large may include non-movable objects); ii) Baselines with manually defined heuristics can create excessive redundant actions. They tend to pick up everything, making exploration inefficient. In fact, our RoboEXP can be viewed as a special form of a frontier-based method with LMM-guided action selection, which demonstrates better efficiency and effectiveness in our interactive exploration task.
C.3.2 Comparison of Efficiency
The number of actions is only comparable when all hidden objects are found, all unexplored spaces are fully explored, and all object states are recovered after exploration. Given that the Random and Heuristic-Open baselines do not provide comparable exploration results, we primarily compare the efficiency of Heuristic-Full, GPT-4V, and our RoboEXP system. We report the average number of actions in settings where all methods find all hidden objects, fully explore the unexplored space, and restore the object states after exploration in the above table in Tab. 3. The results show that our RoboEXP consistently takes fewer actions to construct the full ACSG during the interactive exploration process and is more aligned with human action choices.
C.4 Extreme Illumination and Random Background
RoboEXP is robust to extreme lighting conditions and complex backgrounds. To demonstrate this, we tested under four different scenarios, each with varying lighting conditions and random backgrounds. Fig. 10 shows twenty different settings and their corresponding results. In various conditions and scenarios, our system is able to successfully conduct interactive exploration and construct the ACSG, indicating the robustness of RoboEXP to these factors.
Table 4: Quantitative Results on Different LMMs. We conduct experiments with GPT-4V and LLaVA acting as the core of the RoboEXP decision module, under the same fifteen settings as in Fig. 10.
[…]
Ours (LLaVA) 25$\pm$25.6 50$\pm$29.6 100$\pm$0.0 23$\pm$21.9 2.5$\pm$0.98
Ours (GPT-4V) 95$\pm$12.9 98$\pm$4.3 100$\pm$0.0 0$\pm$0.0 0.1$\pm$0.26
C.5 Performance on Different LMMs
RoboEXP is compatible with different multimodal foundation models beyond GPT-4V. We conducted additional experiments using the latest LLaVA-v1.6-34b as the core of our decision module and compared it against GPT-4V under the same settings. Tab. 4 shows that both models can work with our RoboEXP system, yet the capacity of LMMs does influence the overall performance. In general, GPT-4V achieves a higher success rate and more consistent behaviors across different scenarios.
C.6 Human Intervention
@@
Our RoboEXP system possesses the capability to autonomously adapt to changes in the environment. We employ two types of human interventions to demonstrate these points (refer to App. D).
@@
C.7 Room-Level Household Scenarios
RoboEXP can work well in room-level household environments. To demonstrate this, we conducted two experiments within an apartment (see Fig. 12), specifically in the dining area and bedroom. We integrated four RGB-D observations captured by a handheld RealSense D455 with ICP-based multi-way alignment. Our system successfully constructs corresponding scene graphs within the room-level household environments. Once the static scene graph is constructed, our decision module effectively identifies the correct objects for exploration. Specifically, it accurately identifies the fridge in Fig. 12 (a) and the cabinet in Fig. 12 (b) for further exploration. Tab. 5 shows the complete responses from GPT-4V in our decision module on determining the actions to take in our two household scenarios.
Figure 12: Experiments on Room-Level Household Scenarios. We conduct experiments in two room-level household environments. The dining area includes a table, fridge, and items on the table, whereas the bedroom includes a bed, cabinet, and a TV. The figure presents (a) the observations with segmentation masks; (b) the low-level reconstructed geometry; (c) the built high-level scene graph.
Appendix D Video Timeline
Scenario A. Exploration-Exploitation
Exploration: 00:43 - 01:16
Exploitation: 01:17 - 01:37
Scenario B. Recursive Reasoning
Exploration: 01:49 - 02:26 (Two scenarios)
Scenario C. Obstruction
Exploration: 02:33 - 02:59
Scenario D. Intervention
Exploration: 03:05 - 04:09 (Two scenarios)
Extension to Mobile Robot
Exploration: 04:13 - 04:53
Table 5: Decision Module Logs for Household Scenarios. We present all responses from GPT-4V for our two household scenarios: the dining space case and the bedroom case. The inputs to our decision module are images from various viewpoints of the query object in a green bounding box.
Query Object LMM Response
Dining Space
Fridge [Analysis]: The object in question is a fridge, which is typically used to store food and drinks. The green bounding box in the images highlights the fridge, which has doors that can be opened. Since the goal is to identify all objects hidden from the current observation in the tabletop setting, and a fridge is an appliance that commonly contains items inside, it is reasonable to assume that opening the fridge could reveal hidden objects. [Final Answer]: Open the doors or drawers.
Chair [Analysis]: The object in question is a chair, which is typically not used to store objects in a way that would require opening or removing parts to reveal them. Chairs are designed for sitting, and while it is possible for items to be placed on or under them, the chair itself does not have compartments that would be opened or closed. The images show the chair from different angles, and there is no indication that there are hidden compartments or objects on the chair that would necessitate an action to uncover them. The chair does not have a top object that would be opened or picked up, and it does not have doors or drawers. [Final Answer]: No action.
Table [Analysis]: The object in question is a table, which is highlighted with a green bounding box in the images provided. From the different viewpoints, it is clear that the table has a flat top surface with no visible drawers or doors that could be opened to reveal hidden objects. The table does not appear to have a covering that could hide something underneath it. The table’s design is simple, with a single central support and no storage compartments. Therefore, there is no action required that would potentially reveal hidden objects in this case. [Final Answer]: No action.
Laptop [Analysis]: The object in question is a laptop, which is placed on a table. Laptops do not have a hidden compartment that would typically contain other objects, nor do they obstruct the view of objects underneath them when they are in an open position, as seen in the images. The laptop is already open, and there is no indication that there is anything on top of it or that it is covering anything beneath it. Therefore, there is no action required that would reveal any hidden objects in this case. [Final Answer]: No action.
Mug [Analysis]: The mug in the green bounding box is sitting on a tabletop and does not appear to be covering anything, nor does it have a lid that could be hiding something inside. Mugs are typically open containers and do not have compartments that would conceal objects. Therefore, there is no need to pick it up to look underneath. [Final Answer]: No action.
Kettle [Analysis]: The object in question is a kettle, which is sitting on top of a table. Kettles are typically not used to store other items, especially when they are placed in a setting like a tabletop where they are likely used for their intended purpose of boiling water. The kettle does not have drawers or doors, and it is unlikely that it is obscuring any items on the table due to its shape and size. Therefore, there is no need to open any part of the kettle, and it is not obstructing any items underneath it. [Final Answer]: No action.
Bedroom
Cabinet [Analysis]: The object in question is a cabinet, as indicated by the green bounding box in the provided images. Cabinets typically have doors or drawers that can be opened to reveal contents stored inside. In the images, the cabinet appears to have several drawers. Since the primary goal is to select an action that has the potential to reveal hidden objects, and the cabinet has drawers that could contain items, the appropriate action would be to open these drawers. [Final Answer]: Open the doors or drawers.
Bed [Analysis]: The object in question is a bed, which is typically not a storage furniture piece with doors or drawers. However, some beds do come with storage options underneath, such as drawers or a lifting mechanism for accessing a storage space. In the provided image, the bed appears to have a solid base without visible drawers or doors. There is no indication that the bed has a mechanism to lift the mattress to reveal storage underneath. Therefore, it is unlikely that there are hidden objects within the bed that can be accessed by opening doors or drawers. [Final Answer]: No action.
Television [Analysis]: The object in question is a television, which is placed on top of a surface, likely a media console or a dresser with drawers. Televisions are not typically used to store objects behind or within them, as they are electronic devices designed for display purposes. The images show the television from two different angles, and there is no indication that there are any hidden compartments or objects behind the screen that would be revealed by moving it. Additionally, televisions are generally not designed to have openable parts that would conceal objects. The media console or dresser beneath the television does have drawers, but since the query is specifically about the television, these are not the focus for the action to be taken. [Final Answer]: No action.